OpenStudy (anonymous):

1) what is the slope of the line that passes through the points (2,7) and (-1,3) a. 2/7 b. 3/4 c. 4/3 d. 1/3

OpenStudy (azureilai):

You can graph it, or just count. 7-3=4 2-(-1) = 3 The slope will be 4/3

OpenStudy (austinl):

\(m=\dfrac{y_2-y_1}{x_2-x_1}\) Where, \((x_1,y_1)\rightarrow~(2,7)\\ (x_2,y_2)\rightarrow~(-1,3)\)